OpenStudy (anonymous):

Need help with radicals! Assuming x > 0, what is the simplified fraction form for 15x√3 / √75x^2 <--all under radiacal I have to type the answer into a box and before it, its says: Numerical Answers Expected! After I multiplied the bottom radical by the top and bottom, I had: 15x√3 * √75x = 15√225x^3 √75x^2 * √75x^2 = 775x^2 Now I have: 15x√225x^3 / 75x^2 Simplified: 225x^2√x / 75x^2 Now I'm not sure where to go from here!

hartnn (hartnn):

is the original question, \(\large \dfrac{15x\sqrt3}{\sqrt{75x^2}}\) ?

OpenStudy (anonymous):

yes, that's how it's supposed to be

hartnn (hartnn):

ok, there was no need to multiply and divide by \sqrt 75 x or \sqrt 75x^2

hartnn (hartnn):

split up 75 into its factors, can you ?

hartnn (hartnn):

\(\sqrt {75}= \sqrt {...\times ...\times ...}\)

OpenStudy (anonymous):

5 and 15?

hartnn (hartnn):

yeah, but 15 is again split able.

OpenStudy (anonymous):

5 * 5 * 3?

OpenStudy (anonymous):

and then 5 is a pair so 5√3 right?

hartnn (hartnn):

yes, now since its a "square" root, you group two numbers and can bring it out of sqrt sign, \(\sqrt {75}= \sqrt {5\times 5\times 3}=5\sqrt3\) yes! you got it even before i started explaning :D

hartnn (hartnn):

so, 15√3 / 5√3 =... ? we will deal with 'x' separately

OpenStudy (anonymous):

3√3 im not sure if it simplifys to 3?

hartnn (hartnn):

yes, its 3 √3 gets cancelled from numerator and denominator , and 15/5 is just 3

OpenStudy (anonymous):

if it is just a plain 3, how do we factor in the x?

hartnn (hartnn):

we will go for 'x' now. numerator =x denominator = \(\sqrt{x^2}\) can you simplify the denominator ?

hartnn (hartnn):

i know you can, you have done a similar thing just a few seconds ago...

OpenStudy (anonymous):

i think you would either simplify to x or maybe just leave it at √x^2?

hartnn (hartnn):

one thing to note \(\sqrt{75x^2}=\sqrt {75}\sqrt{x^2}\) be sure! would it simplify to x ?

OpenStudy (anonymous):

oh, so it would just stay the same right? because it cant be factored?

terenzreignz (terenzreignz):

Medal for "split-able" @hartnn :) @haileemackk some people might be nitpicky... but for instance, what is \[\huge \sqrt{(-5)^2}\] ?

hartnn (hartnn):

lol, terenz... yeah, if you are in doubt, consider 'x' as any number and see what it simplifies to...

OpenStudy (anonymous):

Oh, so √75x^2 would be like saying: √75 * x * x x√75 ?

hartnn (hartnn):

yes! which is 5x√3 as already discussed. so, anything gets cancelled from numerator and denominator ?

terenzreignz (terenzreignz):

My turn to be nitpicky... but isn't \[\huge \sqrt{x^2}=|x|\]?

hartnn (hartnn):

yes, but i don't think it should be taken into consideration here, lets keep it simple :)

terenzreignz (terenzreignz):

Agreed. :)

hartnn (hartnn):

so,\(\sqrt{x^2}=x\)

OpenStudy (anonymous):

i dont understand what you mean about anything being cancelled. wouldn't there have to be two of the same thing to cancel each other out?

hartnn (hartnn):

ok, let me write steps for you.

hartnn (hartnn):

\(\large \dfrac{15x \sqrt 3}{\sqrt {75x^2}}=\dfrac{15x \sqrt 3}{5x \sqrt3}\) clear ?

OpenStudy (anonymous):

would you cancel 15x√3 out?

terenzreignz (terenzreignz):

First, cancel out what can clearly be cancelled out... things that are exactly the same in the numerator and denominator.

hartnn (hartnn):

\(\dfrac{15x \sqrt 3}{\sqrt {75x^2}}=\dfrac{3 \times \cancel{5} \cancel{x} \cancel{\sqrt 3}}{\cancel{5}\cancel{x} \cancel{\sqrt3}}\)

OpenStudy (anonymous):

Ohhh! okay, so now all we have left is the 3?

hartnn (hartnn):

thats it! everything breaks down to just 3 !

terenzreignz (terenzreignz):

Awesome :)

OpenStudy (anonymous):

okay, thank you guys so much for all of your help!

hartnn (hartnn):

you are most welcome ^_^

terenzreignz (terenzreignz):

No problem... though I didn't do much except for minor commentary :)